Problem 5. Let E1 = Q(\sqrt2,\sqrt7 ), E2= (\sqrt2,\sqrt[3]{2}), \alpha 1 = 2\sqrt2 + 7\sqrt7, and \alpha 2 = 2\sqrt2 + 3(\sqrt[3]{2})

(i) Determine [Ei : Q] for i = 1, 2.
(ii) Determine a basis of Ei over Q for i = 1, 2.
(iii) Determine the minimal polynomial of \alpha i over Q for i = 1, 2.
(iv) Determine if each of the extensions E1 / Q and E2 / Q is Galois.
Answer #1

(i). Minimal polynomial of \sqrt{2} over Q is 2 and so (Q(V2): Q) = 2 . Similarly minimal polynomial of \sqrt{7} over Q(V2) is 2 _7 and so (Q(V2, 17): Q(V2)) = 2 . Hence, (E1 : 0) = (Q(V2, 17): Q(V2))(QV2) : 0) = 2 x 2 = 4 .

Minimal polynomial of \sqrt[3]{2} over \mathbb{Q} is とで and so (Q(V2, 72), Q(x2)) = 3. Hence, (E2 : 0) = (Q(V2, 12): Q(12))(Q(12) : ) = 3 x 2 = 6 .

(ii). A basis of Q(V2) over \mathbb{Q} is {1, 2} and a basis of Q[/2, 17) over Q(V2) is LAT} . Hence, a basis of E_1 over \mathbb{Q} is {1, V2, V7, V14) .

A basis of Q[/2, 12) over Q(V2) is {1, 72, 74 . Hence a basis of over \mathbb{Q} is {1, V2, 12, 1272, 74, 72/4 .

(iii). Let a = 2V2 and f(x) = x2 – 343 . Now, f(x – 2/2) = (x – 272)2 - 7 = x2 – 4/2x+8 – 343 = x2 – 4/22 – 335 . And f(x+2V2) = (2+2V2)2 – 7 = 22 +4V2c+8343 = + 4V2c - 335 . So,f (x – 2/2)f(x + 2/2) = (x2 – 4V2.c – 335)(x2 + 4/2x – 335) = (x2 – 335)2 – (472c)2 = x4 – 670.22 +112225 – 32.c. Since 7\sqrt{7} is a root of f(x) , 2V2+7V7 is a root of f(x - 212) and hence the minimal polynomial of 2V2+7V7 is the above 4 degree polynomial.

Take g(2) = 73 – 54 . Then, g(x - 2V2) = (1 - 2V2)3 – 54 = 73 - 6V2.1? + 241 – 16V2 - 54 and 9(2+2/2) = 2.3 +67222 +24x + 16V2 – 54 . Similarly as above, since 3\sqrt[3]{2} is a root of g(x) minimal polynomial of 2024332 is g(x2 - 2V2)9(2+2V2) .

(iv). Note that, E_1 is the splitting field of the polynomial (22 - 2) (1 - 7) over \mathbb{Q} . Since a splitting field is always a normal extension, we get that E_1/\mathbb{Q} is a finite normal extension. Also since characteristics of \mathbb{Q} is 0, any field extension of \mathbb{Q} is separable. Hence, E_1/\mathbb{Q} is a finite, normal, separable extension and so it is a Galois extension.

Now note that, とで is an irreducible polynomial over \mathbb{Q} and E_2 contains a root ( namely, \sqrt[3]{2} ) of this polynomial. Now if this was a normal extension, then E_2 have to contain all the roots the above irreducible polynomial. But \sqrt[3]{2}\omega is a root of the polynomial ( \omega is a complex cube root of 1) and we can clearly see that E_2 does not contain \sqrt[3]{2}\omega . Hence, E2/Q is not a normal extension and so it's not a Galois extension.